OpenStudy (song_of_the_sole):

solve the system of inequality y=-x^2+2x-3 y=4x-3

jimthompson5910 (jim_thompson5910):

Start with the first equation as shown in step 1. In step 2, you'll replace 'y' with '4x-3'. From then on, you get everything to one side and combine like terms. \[\Large y=-x^2+2x-3\] \[\Large 4x-3=-x^2+2x-3\] \[\Large 4x-3+x^2-2x+3=0\] \[\Large x^2+2x=0\] I'll let you finish

OpenStudy (song_of_the_sole):

wait so what do i do

jimthompson5910 (jim_thompson5910):

hint: factor

OpenStudy (song_of_the_sole):

im so confused

jimthompson5910 (jim_thompson5910):

try to factor x^2+2x

OpenStudy (song_of_the_sole):

how do i do that

jimthompson5910 (jim_thompson5910):

what is the common factor of x^2 and 2x?

OpenStudy (song_of_the_sole):

2?

OpenStudy (song_of_the_sole):

am i right?

jimthompson5910 (jim_thompson5910):

think of x^2 as x*x

jimthompson5910 (jim_thompson5910):

what do x*x and 2*x have in common?

OpenStudy (song_of_the_sole):

so x

OpenStudy (song_of_the_sole):

right?

jimthompson5910 (jim_thompson5910):

yes factor out x to get what?

OpenStudy (song_of_the_sole):

1?

jimthompson5910 (jim_thompson5910):

x*x+2*x = x*(....) what goes in the parenthesis?

OpenStudy (song_of_the_sole):

1

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

think of distribution but in reverse

OpenStudy (song_of_the_sole):

wait what?

jimthompson5910 (jim_thompson5910):

for example x*(x+3) = x*x + 3*x = x^2+3x so, x^2+3x = x*(x+3)

OpenStudy (song_of_the_sole):

okay now what

jimthompson5910 (jim_thompson5910):

so what does x^2+2x factor to?

OpenStudy (song_of_the_sole):

im not sure

jimthompson5910 (jim_thompson5910):

use the example I posted

OpenStudy (song_of_the_sole):

i still dont get it this is all so confusing to me im sorry

jimthompson5910 (jim_thompson5910):

ok one moment

jimthompson5910 (jim_thompson5910):

x times what will get you x^2?

OpenStudy (song_of_the_sole):

x

jimthompson5910 (jim_thompson5910):

x times what will give 2x?

OpenStudy (song_of_the_sole):

2

jimthompson5910 (jim_thompson5910):

so hopefully you can see how x^2+2x = x*(x+2) ??

OpenStudy (song_of_the_sole):

ok...?

jimthompson5910 (jim_thompson5910):

if x*(x+2) = 0 then x = 0 or x+2 = 0 ----> x = ???

OpenStudy (song_of_the_sole):

2?

jimthompson5910 (jim_thompson5910):

x+2 = 0 leads to x = ??? it's not x = 2

OpenStudy (song_of_the_sole):

1

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

SOMETHING plus 2 = 0 what is that SOMETHING

OpenStudy (song_of_the_sole):

0

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try to isolate x

OpenStudy (song_of_the_sole):

this is making no sense i dont get it what can be added to 2 to make 0??

jimthompson5910 (jim_thompson5910):

we have x+2 = 0 what must happen to both sides to move the 2 over?

OpenStudy (song_of_the_sole):

subtract

jimthompson5910 (jim_thompson5910):

subtract what

OpenStudy (song_of_the_sole):

2

jimthompson5910 (jim_thompson5910):

subtract 2 from both sides and you get what?

OpenStudy (song_of_the_sole):

-2

OpenStudy (song_of_the_sole):

ohhh -2+2=0

OpenStudy (song_of_the_sole):

right?

jimthompson5910 (jim_thompson5910):

yes so either x = 0 or x = -2

OpenStudy (song_of_the_sole):

what about y?

jimthompson5910 (jim_thompson5910):

if x = 0 then y=4x-3 = ???

OpenStudy (song_of_the_sole):

y=4(0)-3

OpenStudy (song_of_the_sole):

y=-3?

jimthompson5910 (jim_thompson5910):

yes so (0,-3) is one ordered pair solution. There is one more ordered pair solution

jimthompson5910 (jim_thompson5910):

if x = -2, then y=4x-3 = ???

OpenStudy (song_of_the_sole):

y=4(-2)-3 -8-3 -11 y=-11

jimthompson5910 (jim_thompson5910):

yep so the two solutions are `(0,-3)` and `(-2,-11)`

jimthompson5910 (jim_thompson5910):

If you graph the two equations, the two curves will cross at (0,-3) and (-2,-11)

OpenStudy (song_of_the_sole):

i dont need to graph but thank you so much!!

jimthompson5910 (jim_thompson5910):

no problem